### v1

@@ Abstract
Deep Neural Networks (DNN) are crucial in approximating nonlinear functions across diverse applications, ranging from image classification to control. Verifying specific input-output properties can be a highly challenging task due to the lack of a single, self-contained framework that allows a complete range of verification types. To this end, we present ModelVerification.jl (MV), the first comprehensive, cutting-edge toolbox that contains a suite of state-of-the-art methods for verifying different types of DNNs and safety specifications. This versatile toolbox is designed to empower developers and machine learning practitioners with robust tools for verifying and ensuring the trustworthiness of their DNN models.
@@
Deep Neural Network Verification Toolbox Adversarial Robustness
@@ 1 Introduction
The use of Deep Neural Networks (DNNs) is becoming increasingly prominent in several applications, including image classification [12, 21, 22], autonomous navigation [6, 31, 37], robotics [1, 16, 32], and control [25, 41, 42, 43]. The main characteristic of these functions is their ability to approximate complex nonlinear functions often employed in solving these tasks. Nonetheless, while these functions are very efficient, their opaque nature can result in unpredictable and potentially unsafe behavior when small changes in the input, often imperceptible to the human, are performed. Broadly speaking, these functions are subject to so-called “adversarial inputs” [36] that can make them behave unsafely both for the system itself and especially for those around them. Hence, given the applications of DNNs in safety-critical contexts where human life is potentially at risk, the need to obtain formal guarantees about the safety of these systems arises and is of paramount importance.
@@
Figure 1: The user provides the network, the safety property to check, and the desired solver. ModelVerification.jl provides an assertion of whether the safety property holds.
@@ 2 Toolbox Features
Addressing the comprehensiveness, our toolbox provides the ability to perform several types of verification (Fig. 2), not only safety and robustness, using reachability analysis (Fig. 2a-b), but also the possibility to perform adversarial attack (Fig. 2c) –by exploiting one of the main methods, such as Fast Gradient Sign Method (FGSM), Projected Gradient Descent (PGD) attack, and Auto-PGD. Moreover, our toolbox includes recent exact and approximation methods [46, 29, 30] even to enumerate the set of (un)safe regions of a given safety property (Fig. 2d). Finally, ModelVerification.jl provides the user with visual representations of the intermediate results of the verification process (i.e., the reachable sets) as depicted in Fig. 3.
We also introduce a new type of input set, ImageConvexHull, which contains all possible interpolations of the given seed images. ImageConvexHull is particularly useful for semantic perturbations such as occlusion, rain, fog, and shadow.
@@
In the literature, it is worth noting that some solvers work best exploiting GPU computation, while others heavily rely on the CPU. Crucially, our toolbox supports both GPU and CPU-based methods. This dual support, in combination with an elegant, highly modularized, and well-documented BaB design, enables a key feature of our toolbox with respect to other state-of-the-art methods, as such, the possibility to combine different solvers for any verification purpose. We report in Fig. 4 a high-level overview of the computational flow of ModelVerification.jl. As discussed, each base submethod that composes the verify function is highly customizable based on the user’s necessity.
4) Efficiency. Besides prioritizing user-friendliness, the last main feature of ModelVerification.jl is concerned with efficiency. Our toolbox provides significant improvements over the first Julia toolbox ever for FV of DNNs called NeuralVerification.jl (NV.jl) [23]. In detail, NV.jl is written in Julia to provide the community with pedagogical and immediate-to-understand implementations, similar to our goal. However, given the pedagogical nature adopted, performance is suboptimal. In contrast, based on the architectural design choices for our toolbox, we are able to achieve user-friendly implementations and, at the same time, efficient results –in terms of verification time and scalability– comparable to, or in certain cases, surpassing those achieved by state-of-the-art solvers, as shown in Sec. 3.
@@ 3.1 Empirical evaluation on VNN benchmarks
We then evaluate ModelVerification.jl on a subset of benchmarks from VNN-COMP’23 [33], ACAS Xu property $\phi_{1}$ for $45$ different networks that have 13K parameters. We compare with the toolboxes that won the first three places: $\alpha$-$\beta$-crown, Marabou, and PyRat. We run our toolbox on an AWS m5 instance following the same VNN-COMP setup [33] as other toolboxes. The results of other toolboxes are directly from the competition. Detailed setting can be found in our toolbox repository. As shown in fig. 6, our toolbox is faster than other toolboxes for most instances of this property, showcasing the efficiency of our toolbox.
@@ Acknowledgements.
This work is in part supported by Boeing and in part supported by mobility grants for non-EU destinations of the University of Verona’s Doctoral School. Tianhao leads the design, implementation and evaluation of the toolbox. Luca and Tianhao lead the paper writing. Kai and Luca lead the documentation and tutorial. Hanjiang and Peizhi contribute to the toolbox implementation and evaluation.

### v2

@@ Abstract
Deep Neural Networks (DNN) are crucial in approximating nonlinear functions across diverse applications, ranging from image classification to control. Verifying specific input-output properties can be a highly challenging task due to the lack of a single, self-contained framework that allows a complete range of various model architecture and input-output properties. To this end, we present ModelVerification.jl (MV.jl) [^1] https://github.com/intelligent-control-lab/ModelVerification.jl, the first comprehensive, cutting-edge toolbox that contains a suite of state-of-the-art methods for verifying different types of DNNs and input-output specifications. This versatile toolbox is designed to empower developers and machine learning practitioners with robust tools for verifying and ensuring the trustworthiness of their DNN models.  [^2] Accepted by CAV 2025.
@@
Deep Neural Network Verification Adversarial Robustness
@@ 1 Introduction
The use of Deep Neural Networks (DNNs) is becoming increasingly prominent in several applications, including image classification [13, 24, 25, 16], autonomous navigation [6, 35, 41], robotics [1, 18, 36, 47], and control [28, 45, 46, 48]. The main characteristic of these functions is their ability to approximate complex nonlinear functions often employed in solving these tasks. Nonetheless, while these functions are very efficient, their opaque nature can result in unpredictable and potentially unsafe behavior when small changes in the input, often imperceptible to the human, are performed. Broadly speaking, these functions are subject to so-called “adversarial inputs” [40] that can make them behave unsafely both for the system itself and especially for those around them. Hence, given the applications of DNNs in safety-critical contexts where human life is potentially at risk, the need to obtain formal guarantees about the safety of these systems arises and is of paramount importance.
@@
Figure 1: The user specifies the network, the safety property to check, and the solver. ModelVerification.jl provides an assertion of whether the safety property holds.
@@ 2 Toolbox Features
Addressing the comprehensiveness, our toolbox provides the ability to perform several types of verification (Fig. 2), not only safety and robustness, using reachability analysis (Fig. 2a-b), but also the possibility to perform adversarial attack (Fig. 2c) –by exploiting one of the main methods, such as Fast Gradient Sign Method (FGSM), Projected Gradient Descent (PGD) attack, and Auto-PGD. Moreover, our toolbox includes recent exact and approximation methods [51, 33, 34] even to enumerate the set of (un)safe regions of a given safety property (Fig. 2d). Finally, ModelVerification.jl provides the user with visual representations of the intermediate results of the verification process (i.e., the reachable sets) as depicted in Fig. 3. We also introduce a new type of input set, ImageConvexHull, which contains all possible interpolations of the given seed images. ImageConvexHull is particularly useful for semantic perturbations such as occlusion, rain, fog, and shadow.
@@
In the literature, it is worth noting that some solvers work best exploiting GPU computation, while others heavily rely on the CPU. Crucially, our toolbox supports both GPU and CPU-based methods. This dual support, in combination with an elegant, highly modularized, and well-documented BaB design, enables a key feature of our toolbox with respect to other state-of-the-art methods, as such, the possibility to combine different solvers for any verification purpose. We report in Fig. 4 a high-level overview of the computational flow of ModelVerification.jl. As discussed, each base submethod that composes the verify function is highly customizable based on the user’s necessity. Based on MV.jl, neural control barrier functions [20] and neural Hamilton-Jacobi Reachability value functions [52] can be verified.
4) Efficiency. Besides prioritizing user-friendliness, the last main feature of ModelVerification.jl is concerned with efficiency. Our toolbox provides significant improvements over the first Julia toolbox ever for FV of DNNs called NeuralVerification.jl (NV.jl) [26]. In detail, NV.jl is written in Julia to provide the community with pedagogical and immediate-to-understand implementations, similar to our goal. However, given the pedagogical nature adopted, performance is suboptimal. In contrast, based on the architectural design choices for our toolbox, we are able to achieve user-friendly implementations and, at the same time, efficient results –in terms of verification time and scalability– comparable to, or in certain cases, surpassing those achieved by state-of-the-art solvers, as shown in Sec. 3. Recently, [31] efficiently verifies semantic perturbations on images using zonotope-based reachability analysis, while [7] verifies robust model predictive control leveraging the efficient CROWN [54] implementation in ModelVerification.jl.
@@ 3.1 Empirical evaluation on VNN benchmarks
We then evaluate ModelVerification.jl on a subset of benchmarks from VNN-COMP’23 [37], ACAS Xu property $\phi_{1}$ for $45$ different networks that have 13K parameters. We compare with the toolboxes that won the first three places: $\alpha$-$\beta$-CROWN, Marabou, and PyRat. We run our toolbox on an AWS m5 instance following the same VNN-COMP setup [37] as other toolboxes. The results of other toolboxes are directly from the competition. Detailed setting can be found in our toolbox repository. As shown in Fig. 6, our toolbox is faster than other toolboxes for most instances of this property, showcasing its efficiency.
@@ Acknowledgements.
This work is in part supported by Boeing and in part supported by mobility grants for non-EU destinations of the University of Verona’s Doctoral School. Tianhao leads the design, implementation and evaluation of the toolbox. Hanjiang, Peizhi, and Xusheng contribute to the toolbox implementation and evaluation. Luca and Tianhao lead the paper writing. Kai and Luca lead the documentation and tutorial.